# Resonant W and Z Boson Production in FSRQ Jets: Implications for Diffuse Neutrino Fluxes

### v2

@@ Abstract
Blazars, particularly Flat Spectrum Radio Quasars (FSRQs), are well-known for their ability to accelerate a substantial population of electrons and positrons, as inferred from multiwavelength radiation observations. Therefore, these astrophysical objects are promising candidates for studying high-energy electron–positron interactions, such as the production of $W^{\pm}$ and $Z$ bosons. In this work, we explore the implications of electron–positron annihilation processes in the jet environments of FSRQs, focusing on the resonant production of electroweak bosons and their potential contribution to the diffuse neutrino flux. By modeling the electron distribution in the jet of the FSRQ 3C 279 during a flaring state, we calculate the reaction rates for $W^{\pm}$ and $Z$ bosons and estimate the resulting diffuse fluxes from the cosmological population of FSRQs. We incorporate the FSRQ luminosity function and its redshift evolution to account for the population distribution across cosmic time, finding that the differential flux contribution exhibits a pronounced peak at redshift $z\sim 1$. While the expected fluxes remain well below the detection thresholds of current neutrino observatories such as IceCube, KM3NeT, or Baikal-GVD, the expected flux from the $Z$ boson production could account for approximately $10^{-3}$ of the total diffuse astrophysical neutrino flux. These results provide a theoretical benchmark for the role of Standard Model electroweak processes in extreme astrophysical environments and emphasize the interplay between particle physics and astrophysics, illustrating that even rare high-energy interactions can leave a subtle but quantifiable imprint on the diffuse astrophysical neutrinos.
@@ 1 Introduction
In 1959, S. L. Glashow had predicted that the cross section for electron antineutrino–electron scattering should be resonantly enhanced due to the production of on-shell $W^{-}$ bosons in the s-channel $\bar{\nu}_{e}e^{-}\rightarrow W^{-}$ [33]. Today, this process is usually referred to as the Glashow resonance. To produce the resonance (with a $W^{-}$ boson mass of approximately $80.3\,\mathrm{GeV}$) through scattering with electrons at rest, the incident (anti)neutrino must have an energy of about $6.3\,\mathrm{PeV}$ [34]. Such high-energy neutrinos could come from cosmic rays, but the corresponding relatively low flux on Earth apparently requires significant observation times even with large-scale detectors like IceCube. For instance, after over a decade of observations, IceCube has detected only a few PeV-scale neutrino events, with one candidate event occurring near the resonance energy at $6.05\pm 0.72\,\mathrm{PeV}$ [35]. Although many studies have explored the implications and potential signatures of detection of the Glashow resonance, definitive experimental confirmation of its existence remains elusive [35, 36].
@@ 2.2 Modeling electron energy distribution in 3C 279
We specifically model the FSRQ 3C 279 during its flaring state on December 20, 2013, which has been extensively studied in previous works [e.g., 18, 12, 40, 30, 13]. Table 1 summarizes the parameters used to fit the observed SED of 3C 279 during this flaring state, as obtained in [13]. The SED modeling is based on synchrotron, SSC, and external Compton emissions from accelerated electrons within the jet blob. Since the observed spectrum covers a limited frequency range, multiple combinations of parameters can produce results consistent with the observations. Consequently, there is an intrinsic level of uncertainty in key parameters such as the acceleration parameter $a$, the turbulent diffusion coefficient $D_{0}$, the Doppler factor $\delta_{D}$, and the photon energy densities in the dust and BLR regions. The parameters $a$ and $D_{0}$ govern acceleration, energy losses, and diffusion within the jet blob, while the Doppler factor and photon energy densities determine radiative cooling [e.g., 30, 31]. Since the electron–positron collision processes considered in this work occur at energies $\sim 100$ GeV, the associated uncertainties in these physical parameters can significantly affect the predicted electron distributions and, subsequently, the corresponding reaction rates in the blazar jet.
@@
3 Electron–positron interactions in blazar jets
@@ 3.1 Resonant W boson production via electron–positron interactions in the jet from 3C 279
We examine the possibility of excitation of the Glashow resonance via electron–positron collisions in the jet blob of AGNs. This idea is inspired by a recent theoretical study demonstrating that such a resonance could occur in the absence of a primary neutrino beam [37]. In this scenario, high-energy electron–positron collisions can produce on-shell $W^{\pm}$ bosons through the $s$-channel subprocess in $e^{+}e^{-}\rightarrow W^{\pm}\,\rho(770)^{\mp}$. While this mechanism has been considered in the context of high-luminosity electron–positron colliders, we evaluate its feasibility in astrophysical environments such as blazars, which are regarded as efficient accelerators of electrons and positrons. It is interesting to note that the proposed process allows one to probe the CP conjugate of the Glashow resonance, providing thus an important test of the neutrino sector of the Standard Model in the resonance region.
While particles are accelerated within the compact jet blob region, the electron–positron interactions can occur throughout the entire blazar jet, as accelerated particles may escape the blob and advect along the jet flow. According to the numerical works for acceleration of pair plasma in the relativistic flow [e.g., 41, 42], it is expected that the number of accelerated positrons would be roughly equal to the number of accelerated electrons. Considering the number density of $\sim 50\,\mathrm{GeV}$ electrons and positrons within the jet blob, the production rate of these particles is estimated as
$\dot{N}_{e^{-}}^{50\,\mathrm{GeV}}\sim\dot{N}_{e^{+}}^{50\,\mathrm{GeV}}\sim\frac{N_{e}(\gamma\approx 9.8\times 10^{4})\,\Delta\gamma}{V_{b}}\,\pi R_{b}^{2}c\sim 3.9\times 10^{42}\,\mathrm{s}^{-1}.$ (3.1)
Assuming a typical blazar jet lifetime of $\tau_{\rm jet}\sim 100\,\mathrm{Myr}$, and a jet volume characterized by a length scale $l_{\rm jet}\sim 100\,\mathrm{pc}$ and transverse size $R_{\rm jet}\sim 0.1\,\mathrm{pc}$ based on observational modeling [e.g., 28], the mean number density over the jet scale becomes
$\bar{n}_{e^{-}}^{50\,\mathrm{GeV}}\sim\bar{n}_{e^{+}}^{50\,\mathrm{GeV}}\sim\frac{\dot{N}_{e^{-}}^{50\,\mathrm{GeV}}\,\tau_{\rm jet}}{\pi R_{\rm jet}^{2}l_{\rm jet}}\sim 140.4\,\mathrm{cm}^{-3}.$ (3.2)
@@
$\Gamma_{e^{-}e^{+}}^{W^{\pm}}\sim n_{e^{-}}^{50\,\mathrm{GeV}}\,n_{e^{+}}^{50\,\mathrm{GeV}}\,\sigma_{e^{-}e^{+}}^{W^{\pm}}\,c\,\pi R_{\rm jet}^{2}l_{\rm jet}\sim 5.6\times 10^{28}\,\mathrm{s}^{-1}.$ (3.3)
@@
$L_{e^{-}e^{+}}^{W^{\pm}}\sim 100\,\mathrm{GeV}\times\Gamma_{e^{-}e^{+}}^{W^{\pm}}\sim 9.0\times 10^{26}\,\mathrm{erg\,s^{-1}},$ (3.4)
which is many orders of magnitude smaller than the jet power (i.e., an order of $10^{45}~{\rm erg~s^{-1}}$). While the total number of electron–positron collisions in a blazar jet may reach $\sim 10^{28}\,\mathrm{s}^{-1}$, corresponding to $\gtrsim 10^{35}$ events $\mathrm{yr}^{-1}$, future leptonic colliders like CEPC [43] and FCC-ee [44] are expected to observe an order of 10 events annually under controlled conditions [37]. This comparison highlights the complementarity between high-energy astrophysical environments, which provide vast natural laboratories for rare electroweak interactions, and precision collider experiments, which offer clean, controllable settings to verify the microphysics of such processes.
@@
$F_{e^{-}e^{+}}^{W^{\pm},\rm obs}\sim\frac{\Gamma_{e^{-}e^{+}}^{W^{\pm}}}{4\pi d_{L}^{2}(1+z)}\sim 3.2\times 10^{-28}\,\mathrm{cm^{-2}\,s^{-1}},$ (3.5)
which lies many orders of magnitude below the sensitivity thresholds of current neutrino telescopes such as IceCube [e.g., 45], KM3NET [e.g., 46], and Baikal-GVD [e.g., 47]. We have assumed that each reaction eventually produces a neutrino. This is justified since the $W^{\pm}$ boson predominantly decays either directly into a lepton pair (as $\nu_{e}e^{+}$) or into hadrons, which would subsequently produce neutrinos (through meson decays, for example). Note that, in these decays, neutrinos of all three flavors will appear. Although the Glashow resonance via electron–positron collisions is theoretically possible in astrophysical environments, the extremely low flux anticipated from AGNs renders direct detection virtually unfeasible with existing or near-future instrumentation. Nevertheless, the existence of this interaction channel in high-density astrophysical plasmas remains a theoretically compelling prospect. We have used the conservative value of the cross section. The latter can be modified by contributions beyond the Standard Model as well as by the momentum transfer-dependent couplings of the mesons to the leptonic current [37]. It may inspire future observational strategies or reinterpretations of high-energy astrophysical phenomena, particularly under more extreme or localized conditions.
@@ 3.2 Resonant Z boson production via electron–positron interactions in the jet from 3C 279
$n_{e^{-}}^{50\,\mathrm{GeV}}\sim n_{e^{+}}^{50\,\mathrm{GeV}}\sim 140.4\,\mathrm{cm^{-3}}.$ (3.6)
@@
$\Gamma_{e^{-}e^{+}}^{Z}\sim n_{e^{-}}^{50\,\mathrm{GeV}}\,n_{e^{+}}^{50\,\mathrm{GeV}}\,\sigma_{e^{-}e^{+}}^{Z}\,c\,\pi R_{\rm jet}^{2}l_{\rm jet}\sim 2.1\times 10^{39}\,\mathrm{s^{-1}}.$ (3.7)
@@
$L_{e^{-}e^{+}}^{Z}\sim 100\,\mathrm{GeV}\times\Gamma_{e^{-}e^{+}}^{Z}\sim 3.4\times 10^{37}\,\mathrm{erg\,s^{-1}}.$ (3.8)
@@
$F_{e^{-}e^{+}}^{Z,\rm obs}\sim\frac{\Gamma_{e^{-}e^{+}}^{Z}}{4\pi d_{L}^{2}(1+z)}\sim 1.3\times 10^{-17}\,\mathrm{cm^{-2}\,s^{-1}}.$ (3.9)
This value, while still below the sensitivity of current neutrino telescopes such as IceCube [e.g., 45], KM3NeT [e.g., 46], or Baikal-GVD [e.g., 47], is dramatically enhanced compared to the resonance channel $e^{+}e^{-}\rightarrow W^{\pm}\,\rho(770)^{\pm}$. Importantly, the $Z$ boson decays also produce all three neutrino flavors, implying a broader range of potential astrophysical signatures. Again, due to the dominance of the leptonic and hadronic decay channels, we assume that each $Z$ boson eventually gives a neutrino. Although the predicted flux remains too small for direct detection, the vastly larger cross section of the $Z$ channel suggests that resonant annihilation into $Z$ bosons could play a non-negligible role in shaping the secondary particle environment in dense pair plasmas. As shown in Figure 5, we examine the parameter dependence of $Z$ boson production on acceleration, diffusion, and cooling processes. While the same trends discussed in Section 3.1 are observed, we find that astrophysical uncertainties can lead to variations in the resulting neutrino flux of up to an order of magnitude. This level of uncertainty, however, remains much smaller than the large gap between the expected neutrino flux from 3C 279 and the sensitivity of current neutrino telescopes. Its inclusion alongside the $W^{\pm}$ channel thus provides a more complete picture of leading electroweak processes in blazar jets.
@@
Figure 7: (a) Integrands of $F_{e^{-}e^{+}}^{Z,{\rm diffuse}}$ as functions of luminosity L for different redshifts $z$. At $z>2$, the contribution is dominated by luminous sources with $L>10^{48}{\rm erg~s^{-1}}$. (b) Differential flux contribution, $dF_{e^{-}e^{+}}^{Z,{\rm diffuse}}/dz$ integrated over luminosity, shown for $0.01\leq z\leq 5$. The distribution exhibits a clear peak at $z\sim 1$, reflecting the cosmic evolution of the FSRQ population.

### v3

@@ Abstract
Blazars, particularly Flat Spectrum Radio Quasars (FSRQs), are well-known for their ability to accelerate a substantial population of electrons and positrons, as inferred from multiwavelength radiation observations. Therefore, these astrophysical objects are promising candidates for studying high-energy electron–positron interactions, such as the production of $W^{\pm}$ and $Z$ bosons. In this work, we explore the implications of electron–positron annihilation processes in the jet environments of FSRQs, focusing on the resonant production of electroweak bosons and their potential contribution to the diffuse neutrino flux. By modeling the electron distribution in the jet of the FSRQ 3C 279 during a flaring state, we calculate the reaction rates for $W^{\pm}$ and $Z$ bosons and estimate the resulting diffuse fluxes from the cosmological population of FSRQs. We incorporate the FSRQ luminosity function and its redshift evolution to account for the population distribution across cosmic time, finding that the differential flux contribution exhibits a pronounced peak at redshift $z\sim 1$. While the expected fluxes remain well below the detection thresholds of current neutrino observatories such as IceCube, KM3NeT, or Baikal-GVD, the flux from $Z$ boson production within the jet blob is many orders of magnitude smaller than the total diffuse astrophysical neutrino flux. These results provide a theoretical benchmark for the role of Standard Model electroweak processes in extreme astrophysical environments, highlighting the interplay between particle physics and astrophysics, and illustrating that even extremely rare high-energy interactions can leave a subtle, theoretically meaningful imprint on the diffuse astrophysical neutrino background.
@@ 1 Introduction
In 1959, S. L. Glashow had predicted that the cross section for electron antineutrino–electron scattering should be resonantly enhanced due to the production of on-shell $W^{-}$ bosons in the s-channel $\bar{\nu}_{e}e^{-}\rightarrow W^{-}$ [33]. Today, this process is usually referred to as the Glashow resonance. To produce the resonance (with a $W^{-}$ boson mass of approximately $80.3\,\mathrm{GeV}$) through scattering with electrons at rest, the incident (anti)neutrino must have an energy of about $6.3\,\mathrm{PeV}$ [34]. Such high-energy neutrinos could come from cosmic rays, but the corresponding relatively low flux on Earth apparently requires significant observation times even with large-scale detectors like IceCube. For instance, after over a decade of observations, IceCube has detected only a few PeV-scale neutrino events, with one candidate event occurring near the resonance energy at $6.05\pm 0.72\,\mathrm{PeV}$ [35]. Although many studies have explored the implications and potential signatures of detection of the Glashow resonance, independent experimental confirmations of its existence remain elusive [35, 36].
@@ 2.1 Model description
In this work the stochastic acceleration coefficient $D_{\rm sto}\propto\gamma^{2}$ corresponds to the so called hard sphere scattering regime in which the scattering mean free path is independent of particle energy. Such a prescription is commonly adopted in one zone models of blazar jets and is motivated by numerical studies of strong turbulence with $\delta B/B\gtrsim 1$ [e.g., 18, 19]. In contrast weak turbulence is expected to yield an energy dependent diffusion coefficient $D_{\rm sto}\propto\gamma^{q}$ with $q=1/3$ for Kolmogorov type spectra. To assess the sensitivity to this assumption we have verified that adopting a Kolmogorov like scaling leads to a modest steepening of the steady state electron distribution and results in changes of less than a factor of two in the predicted neutrino flux. Our main conclusions are therefore robust against the specific choice of the stochastic acceleration regime.
We also note that recent studies of particle transport in strongly magnetized and intermittent turbulence have demonstrated that spatial diffusion can exhibit a non trivial energy dependence driven by field line reversals and regions of large magnetic field curvature rather than following a simple Bohm like scaling [e.g., 40]. In the present work, however, our treatment of particle transport is phenomenological and anchored to one zone spectral energy distribution modeling where the electron/positron distribution is constrained by observations. Since the electroweak interaction rates considered here depend primarily on the total pair density in the emitting region and on the fraction of particles near the resonant energy the detailed microphysics of spatial diffusion does not affect our main results at the order-of-magnitude level.
@@ 2.2 Modeling electron energy distribution in 3C 279
We specifically model the FSRQ 3C 279 during its flaring state on December 20, 2013, which has been extensively studied in previous works [e.g., 18, 12, 41, 30, 13]. Table 1 summarizes the parameters used to fit the observed SED of 3C 279 during this flaring state, as obtained in [13]. The SED modeling is based on synchrotron, SSC, and external Compton emissions from accelerated electrons within the jet blob. Since the observed spectrum covers a limited frequency range, multiple combinations of parameters can produce results consistent with the observations. Consequently, there is an intrinsic level of uncertainty in key parameters such as the acceleration parameter $a$, the turbulent diffusion coefficient $D_{0}$, the Doppler factor $\delta_{D}$, and the photon energy densities in the dust and BLR regions. The parameters $a$ and $D_{0}$ govern acceleration, energy losses, and diffusion within the jet blob, while the Doppler factor and photon energy densities determine radiative cooling [e.g., 30, 31]. Since the electron–positron collision processes considered in this work occur at energies $\sim 100$ GeV, the associated uncertainties in these physical parameters can significantly affect the predicted electron distributions and, subsequently, the corresponding reaction rates in the blazar jet.
@@
3 Electron–positron interactions and neutrino emission in blazar jets
In this section, we present the neutrino emission in blazar jets associated with electron–positron interactions. Before introducing the individual interaction channels we clarify the formulation adopted throughout this section. The electroweak bosons $W^{\pm}$ and $Z$ produced in electron–positron collisions are extremely short lived and decay on timescales that are many orders of magnitude shorter than any relevant dynamical timescale of the jet emitting region. Consequently, treating the interaction in terms of resonant on shell $W^{\pm}$ or $Z$ production followed by prompt decay is fully equivalent to directly computing neutrino production from electron–positron annihilation into neutrino final states. We implicitly account for the relevant branching ratios through the decay of the electroweak bosons such that the effective neutrino yield is identical to that obtained from a direct $e^{+}e^{-}\to\nu\bar{\nu}$ treatment. We nevertheless formulate the calculation in terms of explicit $W^{\pm}$ and $Z$ production in order to highlight the role of electroweak resonances their characteristic energy scales and their branching into neutrino final states. This choice provides a transparent physical interpretation of the interaction channels and does not affect the resulting neutrino spectra or flux normalization.
@@ 3.1 Resonant W boson production via electron–positron interactions in the jet from 3C 279
We primarily examine the possibility of excitation of the Glashow resonance via electron–positron collisions in the jet blob of AGNs. This idea is inspired by a recent theoretical study demonstrating that such a resonance could occur in the absence of a primary neutrino beam [37]. In this scenario, high-energy electron–positron collisions can produce on-shell $W^{\pm}$ bosons through the $s$-channel subprocess in $e^{+}e^{-}\rightarrow W^{\pm}\,\rho(770)^{\mp}$. While this mechanism has been considered in the context of high-luminosity electron–positron colliders, we evaluate its feasibility in astrophysical environments such as blazars, which are regarded as efficient accelerators of electrons and positrons. It is interesting to note that the proposed process allows one to probe the CP conjugate of the Glashow resonance, providing thus an important test of the neutrino sector of the Standard Model in the resonance region.
While particles are accelerated within the compact jet blob region, we consider that the dominant electron–positron interactions occur within the blob itself, where the lepton density is the highest. According to numerical works on acceleration of pair plasma in relativistic flows [e.g., 42, 43], the number of accelerated positrons is expected to be roughly equal to the number of accelerated electrons. Using the number density of $\sim 50\,\mathrm{GeV}$ electrons and positrons within the blob, the volume-averaged number density of these particles is estimated as
$\bar{n}_{e^{-}}^{50\,\mathrm{GeV}}\sim\bar{n}_{e^{+}}^{50\,\mathrm{GeV}}\sim\frac{N_{e}(\gamma\approx 9.8\times 10^{4})\,\Delta\gamma}{V_{b}}\,\sim 7.2\,\mathrm{cm}^{-3}.$ (3.1)
@@
$\Gamma_{e^{-}e^{+}}^{W^{\pm}}\sim n_{e^{-}}^{50\,\mathrm{GeV}}\,n_{e^{+}}^{50\,\mathrm{GeV}}\,\sigma_{e^{-}e^{+}}^{W^{\pm}}\,c\,V_{b}\sim 2.9\times 10^{16}\,\mathrm{s}^{-1}.$ (3.2)
We note that the diffusion timescale for electrons to escape the blob is much longer than the typical timescale for electron–electron and electron–positron interactions within the blob. Using the parameters $R_{b}=2.4\times 10^{15}~\mathrm{cm}$, $B=0.21~\mathrm{G}$, $D_{0}=9.0\times 10^{-6}$, and $\gamma=9.8\times 10^{4}$, the diffusive escape timescale can be expressed in terms of the dimensionless escape parameter $\tau$ as
$t_{\rm esc}=\frac{\tau}{\gamma D_{0}}=\frac{R_{b}^{2}eB}{\gamma m_{e}c^{3}}\sim 2.4\times 10^{11}~{\rm s},$ (3.3)
which is many orders of magnitude longer than both the blob dynamical timescale, $t_{\rm dyn}\sim R_{b}/c\sim 8\times 10^{4}~{\rm s}$. We note that $\Gamma_{e^{-}e^{+}}^{W^{\pm}}$ represents the total interaction rate integrated over the blob volume, corresponding to a characteristic global interaction timescale $t_{\rm e^{-}e^{+}}^{\rm W^{\pm}}\sim 1/\Gamma_{e^{-}e^{+}}^{W^{\pm}}\sim 3\times 10^{-17}~\mathrm{s}$. This ensures that essentially all interactions occur before electrons can escape, justifying the in-blob treatment of electroweak processes.
@@
$L_{e^{-}e^{+}}^{W^{\pm}}\sim 100\,\mathrm{GeV}\times\Gamma_{e^{-}e^{+}}^{W^{\pm}}\sim 4.6\times 10^{15}\,\mathrm{erg\,s^{-1}},$ (3.4)
which is many orders of magnitude smaller than the jet power (i.e., an order of $10^{45}~{\rm erg~s^{-1}}$). While the total number of electron–positron collisions in a blazar jet may reach $\sim 10^{16}\,\mathrm{s}^{-1}$, corresponding to $\gtrsim 10^{23}$ events $\mathrm{yr}^{-1}$, future leptonic colliders like CEPC [44] and FCC-ee [45] are expected to observe an order of 10 events annually under controlled conditions [37]. This comparison highlights the complementarity between high-energy astrophysical environments, which provide vast natural laboratories for rare electroweak interactions, and precision collider experiments, which offer clean, controllable settings to verify the microphysics of such processes.
@@
$F_{e^{-}e^{+}}^{W^{\pm},\rm obs}\sim\frac{\Gamma_{e^{-}e^{+}}^{W^{\pm}}}{4\pi d_{L}^{2}(1+z)}\sim 1.7\times 10^{-40}\,\mathrm{cm^{-2}\,s^{-1}},$ (3.5)
which lies many orders of magnitude below the sensitivity thresholds of current neutrino telescopes such as IceCube [e.g., 46], KM3NET [e.g., 47], and Baikal-GVD [e.g., 48]. We have assumed that each reaction eventually produces a neutrino. This is justified since the $W^{\pm}$ boson predominantly decays either directly into a lepton pair (as $\nu_{e}e^{+}$) or into hadrons, which would subsequently produce neutrinos (through meson decays, for example). Note that, in these decays, neutrinos of all three flavors will appear. Although the Glashow resonance via electron–positron collisions is theoretically possible in astrophysical environments, the extremely low flux anticipated from AGNs renders direct detection virtually unfeasible with existing or near-future instrumentation. Nevertheless, the existence of this interaction channel in high-density astrophysical plasmas remains a theoretically compelling prospect. We have used the conservative value of the cross section. The latter can be modified by contributions beyond the Standard Model as well as by the momentum transfer-dependent couplings of the mesons to the leptonic current [37]. It may inspire future observational strategies or reinterpretations of high-energy astrophysical phenomena, particularly under more extreme or localized conditions.
@@ 3.2 Resonant Z boson production via electron–positron interactions in the jet from 3C 279
$n_{e^{-}}^{50\,\mathrm{GeV}}\sim n_{e^{+}}^{50\,\mathrm{GeV}}\sim 7.2\,\mathrm{cm^{-3}}.$ (3.6)
@@
$\Gamma_{e^{-}e^{+}}^{Z}\sim n_{e^{-}}^{50\,\mathrm{GeV}}\,n_{e^{+}}^{50\,\mathrm{GeV}}\,\sigma_{e^{-}e^{+}}^{Z}\,c\,V_{b}\sim 1.1\times 10^{27}\,\mathrm{s^{-1}}.$ (3.7)
@@
$L_{e^{-}e^{+}}^{Z}\sim 100\,\mathrm{GeV}\times\Gamma_{e^{-}e^{+}}^{Z}\sim 1.8\times 10^{26}\,\mathrm{erg\,s^{-1}}.$ (3.8)
@@
$F_{e^{-}e^{+}}^{Z,\rm obs}\sim\frac{\Gamma_{e^{-}e^{+}}^{Z}}{4\pi d_{L}^{2}(1+z)}\sim 6.8\times 10^{-30}\,\mathrm{cm^{-2}\,s^{-1}}.$ (3.9)
This value, while still below the sensitivity of current neutrino telescopes such as IceCube [e.g., 46], KM3NeT [e.g., 47], or Baikal-GVD [e.g., 48], is dramatically enhanced compared to the resonance channel $e^{+}e^{-}\rightarrow W^{\pm}\,\rho(770)^{\pm}$. Importantly, the $Z$ boson decays also produce all three neutrino flavors, implying a broader range of potential astrophysical signatures. Again, due to the dominance of the leptonic and hadronic decay channels, we assume that each $Z$ boson eventually gives a neutrino. Although the predicted flux remains too small for direct detection, the vastly larger cross section of the $Z$ channel suggests that resonant annihilation into $Z$ bosons could play a non-negligible role in shaping the secondary particle environment in dense pair plasmas. As shown in Figure 5, we examine the parameter dependence of $Z$ boson production on acceleration, diffusion, and cooling processes. While the same trends discussed in Section 3.1 are observed, we find that astrophysical uncertainties can lead to variations in the resulting neutrino flux of up to an order of magnitude. This level of uncertainty, however, remains much smaller than the large gap between the expected neutrino flux from 3C 279 and the sensitivity of current neutrino telescopes. Its inclusion alongside the $W^{\pm}$ channel thus provides a more complete picture of leading electroweak processes in blazar jets.
@@
Figure 7: (a) Integrands of $F_{e^{-}e^{+}}^{Z,{\rm diffuse}}$ as functions of luminosity L for different redshifts $z$. At $z>2$, the contribution is dominated by luminous sources with $L>10^{48}{\rm erg~s^{-1}}$. (b) Differential flux contribution, $dF_{e^{-}e^{+}}^{Z,{\rm diffuse}}/dz$ integrated over luminosity, shown for $0.01\leq z\leq 5$. The distribution exhibits a clear peak at $z\sim 1.5$, reflecting the cosmic evolution of the FSRQ population.